OpenStudy (adamaero):

integrate (x^2)*(2-x)^.5 from 0 to 4

OpenStudy (adamaero):

I got -64/15 or -4.26666

OpenStudy (adamaero):

\[\int\limits_{0}^{4} x^2(2-x)^{1/2} dx\]

OpenStudy (adamaero):

had to do u-sub, right?

OpenStudy (irishboy123):

what do you get if you let x = 4 in the function you are integrating?

OpenStudy (adamaero):

I don't follow. That's what I did, after u-substitution,

OpenStudy (irishboy123):

you are integrating a function, let's call it f(x), where: \[f(x) = x^2 \ (2−x)^\frac{1}{2}\]over the region 0≤x≤4. so you are calculating the area under f(x) over region 0≤x≤4. have you plotted f(x) ?? and what happens to f(x) when x > 2 ??

OpenStudy (adamaero):

No plotting neccissary, it's just an integral.

OpenStudy (adamaero):